OpenStudy (anonymous):

Challenge Part 1 find the area of the region enclosed by the x-axis and the curve y=ln(x^2-x+2) on [0,1]

OpenStudy (anonymous):

Try using integration by parts and later on a partial fraction decomposition.

OpenStudy (anonymous):

\[\large \int1\cdot \ln(x^2-x+2)=x\ln(x^2-x+1)-\int x\cdot \frac{1}{x^2-x+1}(2x-1) \]

OpenStudy (anonymous):

\[\large x\ln(x^2-x+1)-\int\frac{2x^2-x}{x^2-x+1}dx \] Not quite ready for a partial fraction decomposition, but after a longhand division things should look better.

OpenStudy (anonymous):

i should have said, i know this but partial fractions is hard not integration by parts

OpenStudy (fibonaccichick666):

are you sure it is a plus 2 not a minus?

OpenStudy (anonymous):

heck sure...if its minus...its not a challenge anymore

OpenStudy (fibonaccichick666):

if it was a minus, it would be easy

OpenStudy (anonymous):

ya i know

OpenStudy (fibonaccichick666):

did you try integration tables or are you not allowed them?

OpenStudy (anonymous):

If not via partial fraction decomposition, then I can only guess that you want to use a trig identity.

OpenStudy (anonymous):

i dont even know what integral tables are

OpenStudy (fibonaccichick666):

they would solve this in an instant lol

OpenStudy (anonymous):

haha then i guess we are not allowed to use them

OpenStudy (fibonaccichick666):

ready?

OpenStudy (fibonaccichick666):

this is only the \[\int{(x^2-x+2})dx\] I am not carrying the \[2\pi\] around through the whole thing

OpenStudy (fibonaccichick666):

So to start let's look at the polynomial and try to "simplify" it. We can complete the square to get it in the form \[\int{\ln{|(x-.5)^2+1.75|}}dx\]

OpenStudy (fibonaccichick666):

from here let u=x-.5 then du=dx

OpenStudy (fibonaccichick666):

then we have ∫ln|(u)^2+1.75|dx

OpenStudy (fibonaccichick666):

now this is equivalent to \[\int\limits_{}^{} \ln[(u+\frac{\sqrt{7} }{ 2 })(u-\frac{\sqrt{7} }{ 2 })]dx \]

OpenStudy (fibonaccichick666):

that should be a du not dx

OpenStudy (fibonaccichick666):

now by ln rules and int rules we can make it ∫ln[(u+7√2)]du+∫ln[(u−7√2)]du

OpenStudy (fibonaccichick666):

now, yet another substitution let to get ∫ln|v|dv+∫ln|g|dg

OpenStudy (fibonaccichick666):

then plug u back in first then put your x's back

OpenStudy (anonymous):

Careful, it seems to me that you're missing something at one step.

OpenStudy (fibonaccichick666):

final answer that I got in terms of u was was \[\int ln (x^2-x+2) dx=(u+\frac { \sqrt{7}}{2})ln|(u+\frac { \sqrt{7}}{2})|-(u+\frac { \sqrt{7}}{2})+(u-\frac { \sqrt{7}}{2})ln|(u-\frac { \sqrt{7}}{2})+(u-\frac { \sqrt{7}}{2})+C\]

OpenStudy (fibonaccichick666):

Yea, I accidentally dropped the ln at the start but I picked it back up

OpenStudy (anonymous):

You said yourself that u^2+1.75, but you cannot factor this by (u-x)(u+x)=u^2-x^2

OpenStudy (anonymous):

or do I misread that step?

OpenStudy (anonymous):

(Sorry I am on my windows machine, it has problems with LaTeX at some point, so I might have confused it)

OpenStudy (fibonaccichick666):

yea, I used conjugates where 1.75=((sqrt7)/2)^2

OpenStudy (fibonaccichick666):

right or did I screw that step up? It is quite possible

OpenStudy (anonymous):

I see that, but that would resunt in a minus still, not in a plus.

OpenStudy (fibonaccichick666):

ahh it is a minus, alas

OpenStudy (anonymous):

it's really a good idea though, I was having troubles with the plus sign before myself, I tried with getting it into the complex plane first, there it would be possible for this substitution, but the result remains imaginary.

OpenStudy (fibonaccichick666):

yea, I did that first as well. so now at least it is just \[\int(ln|u^2+1.75|du)\]

OpenStudy (fibonaccichick666):

hmm curious

OpenStudy (fibonaccichick666):

if you take the squareroots for it and multiply them together, would trig sub be helpful?

OpenStudy (anonymous):

I dont know why but I like to bring challenge question...so complicated and so interesting...

OpenStudy (fibonaccichick666):

haha yea you do

OpenStudy (anonymous):

the trig substitution looks way more promising than everything else I tried so far, but it remains complex, I can relog reload quick and then try to share what I got.

OpenStudy (fibonaccichick666):

ooh wait why wolfram alpha this

OpenStudy (fibonaccichick666):

why not*

OpenStudy (anonymous):

So far this seems interesting \[\Large \ln\left((x-\frac{1}{2})^2+\frac{7}{4}\right) \] After carrying through with the trig substitution it looks like this: \[\Large \ln\left(\left(\frac{\sqrt{7}}{2}\tan\alpha\right)^2 \right)\cdot \frac{\sqrt{7}}{2}\sec\alpha\tan\alpha \]

OpenStudy (anonymous):

If I didn't make any careless mistakes.

OpenStudy (anonymous):

setting \[\Large u=\frac{\sqrt{7}}{2} \sec\alpha\] to obtain the above.

OpenStudy (fibonaccichick666):

but you still have a ln multiplication

OpenStudy (anonymous):

optically it's pretty neat, especially with the exponent which could also be factored about, but it doesn't help with any substitution I can think of.

OpenStudy (anonymous):

yes, \[\Large 2\ln\left(\left(\frac{\sqrt{7}}{2}\tan\alpha\right) \right)\cdot \frac{\sqrt{7}}{2}\sec\alpha\tan\alpha d\alpha\]

OpenStudy (anonymous):

that's as good as it gets for me, until now.

OpenStudy (fibonaccichick666):

interesting, but i'm not sure how to do that trig sub. but I will def look at it when I have time again

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

interesting stuff...i missed everything lol

OpenStudy (anonymous):

So far it's an useless result anyway, no substitution will help with the above integral, I mean, the one I tried to derive. (-:

OpenStudy (fibonaccichick666):

2.64575 tan^(-1)(0.755929 x)+x (-2.+log(1.75+x^2)) - This is wolfram answer

OpenStudy (fibonaccichick666):

it is wayy tricky

OpenStudy (anonymous):

I understand why there is an arc tan it it, that also speaks at least somehow for a trig substitution, maybe not quite the one I did though.

OpenStudy (anonymous):

check this out

OpenStudy (anonymous):

with parts, \[\Large x\ln(x^2-x+2)-\int\frac{2x^2-x}{x^2-x+2}dx \\ \Large x\ln(x^2-x+2)-\int\left(\frac{x}{x^2-x+2}-\frac{4}{x^2-x+2}+2\right) \]

OpenStudy (anonymous):

this seems promising, is it what they/you did?

OpenStudy (anonymous):

hmm almost. the above could be integrated, will be arctan functions, just messy to deal with.

OpenStudy (anonymous):

\[\Large \int \left(\frac{x}{(x-.5)^2+1.75}+4\frac{1}{(x-.5)^2+1.75}+2\right)dx \]

OpenStudy (anonymous):

so whats the answer...

OpenStudy (anonymous):

well do you agree with the by parts method so far?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[\large \int\ln(x^2-x+2)=x\ln(x^2-x+2)-\int\frac{2x^2-x}{x^2-x+2}dx \] So from here on I will just talk about the very right integral, or more generally, the quotient : \[\Large \frac{2x^2-x}{x^2-x+2}=\frac{2x^2}{x^2-x+2}-\frac{x}{x^2-x+2}\] or after a long division \[\Large \frac{2x^2-x}{x^2-x+2}=2+\frac{2x-4}{x^2-x+2}-\frac{x}{x^2-x+2} \]

OpenStudy (anonymous):

\[\Large \frac{2x^2-x}{x^2-x+2}=2+\frac{x}{x^2-x+2}-\frac{4}{x^2-x+2} \]

OpenStudy (anonymous):

\[\Large \frac{2x^2-x}{x^2-x+2}=2+\frac{x}{(x+\frac{1}{2})^2+\frac{7}{4}}-\frac{4}{x+\frac{1}{2})^2+\frac{7}{4}} \]

OpenStudy (anonymous):

These terms can be integrated.

OpenStudy (anonymous):

The first one will just give a 2x, after that comes a logarithmic property (u sub) with an arctan function and the last one again is an arctan function, so if I didn't make any mistakes, this is in harmony with the result delivered by the wolf. http://www.wolframalpha.com/input/?i=int+ln%28x%5E2-x%2B2%29

OpenStudy (anonymous):

got it...i dont trust wolframalpha for calculus stuff...and check the three pages i posted before...are they correct...those r done by wolframalpha

OpenStudy (anonymous):

Hmm don't quite remember all the steps, I would have to review, but most of it seemed clear to me. It's a good thing to mistrust the wolf in general. The steps by wolf seem a bit lengthy though.

OpenStudy (anonymous):

ya i know

OpenStudy (anonymous):

Do you concur with these steps @FibonacciChick666 ?

OpenStudy (radar):

I think this is above my pay-grade, maybe when I grow up I could help lol.